OpenStudy (rayman5766):

Use transformations to solve the equation. c – 16 = 41 A. c = –57 B. c = –25 C. c = 25 D. c = 57

OpenStudy (anonymous):

add 16 to 41

OpenStudy (carolina4567):

16 + 41 = ?

OpenStudy (rayman5766):

can some one help me I have five questions also I chooson c

OpenStudy (carolina4567):

WHY C what is 16 + 41

OpenStudy (carolina4567):

Its a no brainer

OpenStudy (rayman5766):

sorry i know its 57

OpenStudy (carolina4567):

Ok then ur answer is ?????

OpenStudy (rayman5766):

which is d positive 57

OpenStudy (rayman5766):

What is the solution to the equation? b + 0.8 = –1

OpenStudy (anonymous):

16+41=57